OpenStudy (anonymous):

3x^2+19y+20y^2 factor

OpenStudy (nathan917):

Simplifying 3x2 + 19y + 20y2 = 0 Solving 3x2 + 19y + 20y2 = 0 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-19y' to each side of the equation. 3x2 + 19y + -19y + 20y2 = 0 + -19y Combine like terms: 19y + -19y = 0 3x2 + 0 + 20y2 = 0 + -19y 3x2 + 20y2 = 0 + -19y Remove the zero: 3x2 + 20y2 = -19y Add '-20y2' to each side of the equation. 3x2 + 20y2 + -20y2 = -19y + -20y2 Combine like terms: 20y2 + -20y2 = 0 3x2 + 0 = -19y + -20y2 3x2 = -19y + -20y2 Divide each side by '3'. x2 = -6.333333333y + -6.666666667y2 Simplifying x2 = -6.333333333y + -6.666666667y2 Reorder the terms: x2 + 6.333333333y + 6.666666667y2 = -6.333333333y + 6.333333333y + -6.666666667y2 + 6.666666667y2 Combine like terms: -6.333333333y + 6.333333333y = 0.000000000 x2 + 6.333333333y + 6.666666667y2 = 0.000000000 + -6.666666667y2 + 6.666666667y2 x2 + 6.333333333y + 6.666666667y2 = -6.666666667y2 + 6.666666667y2 Combine like terms: -6.666666667y2 + 6.666666667y2 = 0.000000000 x2 + 6.333333333y + 6.666666667y2 = 0.000000000